How do you find (f - g)(x) for f(x) = x2 + 10x, and g(x) = - 5x + 3 ?
Answer by Theo(13342)

f-g(x) is equivalent to f(x) - g(x)
if f(x) = x^2 + 10x and g(x) = -5x + 3 then:
f-g(x) = (x^2 + 10x) - (-5x + 3)
remove parentheses to get:
f-g(x) = x^2 + 10x + 5x - 3
combine like terms to get:
f-g(x) = x^2 + 15x = 3
here's a reference discussing operations on functions.
f-g(x) is part of that, as is:
f+g(x)
f*g(x)
f/g(x)
fog(x) which is otherwise shown as f(g(x))
in general, these are the rules:
f-g(x) = f(x) - g(x)
f+g(x) = f(x) + g(x)
f*g(x) = f(x) * g(x)
f/g(x) = f(x) / g(x)
fog(x) = f(g(x)) which means f of g of x.
I could go on but the reference says it all so check out the reference if you need more.
http://www.wtamu.edu/academic/anns/mps/math/mathlab/col_algebra/col_alg_tut30b_operations.htm